Another in MAROULIES un y Margales IR BURGER CHEESEBURGER FRENCH PRIES ONION RINGS 14 doing Same here..lin doire is to Sup1) Commentary on the cartoon: The older worker (Boomer) did not plan well for retirement, and has to work to supplement his retirement check. The Social Security system spends money taken from current workers to pay the benefits of today's retirees. That explains the younger worker's (Generation X) response to the older worker. Money is taken from his paycheck to fund the older worker's social security check.

In the near future, there will not be enough people working to pay all retirees at today's rate. The Social Security system will go bankrupt. Social Security's options are to increase the money taken from the workers paychecks or decrease the amount of money paid to retirees. Neither option is popular. An alternative is to establish your own retirement fund. We can use the formulas that we worked with in this Module to analyze different retirement plans.


Your Retirement: Plan 1 - Starting Early ( Please try to answer all the questions if possible)
a) Enter your desired amount of money you wish to receive each month while retired = $ 2000 ( This is my answer)
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT

b) Upon retirement (40 years from now) you want to receive "x dollars" (amount of money you choose in Part a) each month for a period of 28 years. This money resides in an account that pays 5.7% interest. How much money must be in the account to fulfill your goal?
Amount of money needed = $
.
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT. Round your answer to 2 decimal places as needed.

c) You plan to deposit money every month for 40 years into an account that pays 6.1%. What monthly deposit must you make to achieve the amount of money determined in the answer to Part b above?
Payment amount = $
.
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT. Round your answer to 2 decimal places as needed.

d) How much money did you actually deposit into the account?
Total amount deposited = $
.
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT. Round your answer to 2 decimal places as needed.

e) How much money will you actually receive in payments during retirement?
Total amount received = $
.
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT. Round your answer to 2 decimal places as needed.

Your Retirement: Plan 2 - Waiting Until Middle Agef) You wait until later to start saving for retirement, and so are depositing money every month for only 20 years into an the same account that still pays 6.1%. What monthly deposit must you make to achieve the amount of money determined in the answer to Part b above?
Payment amount = $
.
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT. Round your answer to 2 decimal places as needed.

g) How much money did you actually deposit in this plan?
Total amount deposited = $
.
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT. Round your answer to 2 decimal places as needed.

h) How many times more costly is the monthly payment of Plan 2 (saving for 20 years) compared to Plan 1 (saving for 40 years)?
(Hint: Divide Part f by Part c).
How many times more costly 20 year plan is compared to 40 year plan = .
DO NOT TYPE COMMAS OR ANSWERS WILL BE MARKED INCORRECT. Round your answer to 2 decimal places as needed.
Answer #1

a)
2000

b)
=2000/(5.7%/12)*(1-1/(1+5.7%/12)^(12*28))=335379.746604321

c)
=2000/(5.7%/12)*(1-1/(1+5.7%/12)^(12*28))*(6.1%/12)/((1+6.1%/12)^(12*40)-1)=163.890604225588

d)
=2000/(5.7%/12)*(1-1/(1+5.7%/12)^(12*28))*(6.1%/12)/((1+6.1%/12)^(12*40)-1)*12*40=78667.490028282

e)
=2000*12*28=672000


P.S.: I am not allowed to answer more than 4 questions